### v1

# Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions

Abstract.
Explanations of AI systems rarely address the information needs of people affected by algorithmic decision-making (ADM). This gap between conveyed information and information that matters to affected stakeholders can impede understanding and adherence to regulatory frameworks such as the AI Act (Commission, 2021). To address this gap, we present the ”XAI Novice Question Bank”: A catalog of affected stakeholders’ information needs in two ADM use cases (employment prediction and health monitoring), covering the categories data, system context, system usage, and system specifications. Information needs were gathered in an interview study where participants received explanations in response to their inquiries. Participants further reported their understanding and decision confidence, showing that while confidence tended to increase after receiving explanations, participants also met understanding challenges, such as being unable to tell why their understanding felt incomplete. Explanations further influenced participants’ perceptions of the systems’ risks and benefits, which they confirmed or changed depending on the use case. When risks were perceived as high, participants expressed particular interest in explanations about intention, such as why and to what end a system was put in place. With this work, we aim to support the inclusion of affected stakeholders into explainability by contributing an overview of information and challenges relevant to them when deciding on the adoption of ADM systems. We close by summarizing our findings in a list of six key implications that inform the design of future explanations for affected stakeholder audiences.
1. Motivation
Society is in the process of negotiating how AI systems should be used in public institutions and organizations (Züger and Asghari, 2023; Lucaj et al., 2023; Lima et al., 2023). These systems should be transparent, accountable, and support human oversight, according to the legal frameworks which are being developed by regulatory bodies (Commission, 2021). But realizing these general principles in practice can be difficult as they do not always fit specific use cases. Research in the field of explainable AI (XAI) addresses this issue by examining how AI systems can be made understandable to stakeholders and how the given information enables certain actions, such as contesting a decision (Alfrink et al., 2022) or assigning accountability (Langer et al., 2021). As explanations thereby assume a key role in public AI (Züger and Asghari, 2023), the question of what constitutes a good explanation becomes central.
Good explanations of AI systems are aligned with the information needs of the stakeholder audience (Miller, 2019; Mueller et al., 2019; Byrne, 2023; Ehsan et al., 2022) and are a means to help them achieve their aims, which can depend on the stakeholder’s role (Langer et al., 2021; Barredo Arrieta et al., 2020), AI literacy (Long and Magerko, 2020), and domain knowledge (Wang and Yin, 2021). However, information needs of ”affected stakeholders” – people impacted by algorithmic decision-making (ADM) (European Parliament. Directorate General for Parliamentary Research Services., 2019; Golpayegani et al., 2023) – have not yet been explored in-depth in XAI research. This can limit affected stakeholders’ options to understand ADM and, in consequence, infringe on principles of responsible and trustworthy AI (Floridi, 2019; on Artificial Intelligence, 2019; Thiebes et al., 2021; Baeza-Yate and Matthews, 2022). Concretely, providing explanations for affected stakeholders can be essential to reduce the risks of power asymmetry (Ananny and Crawford, 2018), inequality (Lopez, 2019), and informational unfairness (Schoeffer et al., 2022) in deployments of public AI (Kuziemski and Misuraca, 2020; Henman, 2020; Brown et al., 2019; de Fine Licht and de Fine Licht, 2020; Veale et al., 2018).
However, building explanations to increase affected stakeholders’ understanding can be challenging: First, the group of affected stakeholders is diverse and can include ”laypeople” (Shen et al., 2020; Szymanski et al., 2021; Schmude et al., 2023a) or ”AI novices”111We use the term in reference to ”visualization novices” (Burns et al., 2023)., i.e., people with little technical literacy who have heterogeneous information needs, aims, and prior knowledge, necessitating accessible and adaptive explanations (Shulner-Tal et al., 2022; Conati et al., 2021; Kim et al., 2024). Second, the effects of explanations are influenced by perceptions of the deploying institution (Brown et al., 2019; Ehsan et al., 2021), meaning that if the institution is not trusted, the ADM system and its explanations are not trusted either (Woodruff et al., 2018; Scott et al., 2022). Third, understanding is an interconnection of knowledge and reasoning (Grimm, 2019; Keil, 2006) involving numerous cognitive processes (Keil, 2006; Lombrozo and Wilkenfeld, 2019) that can barely be captured by standardized large-scale tests (Sato et al., 2019). And lastly, AI lifecycles encompass large amounts of information (Dhanorkar et al., 2021), necessitating a preselection of what is deemed relevant.
These interactions between cognitive, perceptional, and organizational dimensions can be difficult to address in explanation design. Recent XAI research, therefore, advocates for providing tailored explanations, based on the stakeholder’s aims (Dhanorkar et al., 2021; Freiesleben and König, 2023) and information needs (Liao et al., 2020, 2021), instead of explaining what is assumed to be relevant (Miller, 2023). In line with literature arguing that more research is needed on human relations with AI (Capel and Brereton, 2023), this study examines how to address the information needs of affected stakeholders and meet the challenges outlined. In this paper, we aim to support the inclusion of affected stakeholders into explainability by presenting an empirical study of their information needs, understanding, and perceptions. To this end, we use two ADM use cases as examples: an employment prediction algorithm and a health wristband for elderly care. To allow for the collection of participants’ information needs, the explanation setting is flipped: Participants independently inquire information about the systems to the end of evaluating their adoption, and are provided verbal question-driven explanations in response. We aim to address the challenges outlined above by letting participants choose independently which information they inquire, e.g., about the deploying institution or the algorithm used, reducing the amount of pre-selected information in favor of demand-based information. We analyze how the given information influences participants’ self-reported understanding and decision confidence, and how ”explanatory stances” and challenges in understanding impact participants’ information acquisition. Lastly, we analyze participants’ perceptions of risks and benefits before and after receiving explanations and describe perceived changes. Our work is thus guided by the following research questions:
RQ1-Information Needs: What are affected stakeholders’ information needs when deciding over the adoption of an ADM system?
RQ2-Understanding: How do question-driven explanations interact with participant understanding?
RQ3-Perception: How do question-driven explanations impact perceptions of the systems’ risks and benefits?
The contributions of our work include i) a catalog of affected stakeholders’ questions titled the ”XAI Novice Question Bank” (Figure 2), in reference to Liao et al. (2020)’s XAI Question Bank; ii) recommendations on how to address understanding challenges such as blind spots and outsourcing and on how to use explanatory stances in question-driven explanations; and iii) reflections on the relevance of intentional information in explanations for affected stakeholders and considerations of how to incorporate insights from ”civic education” (Lupia, 2016) into explainability.
2. Related Work
2.1. Human-Centered Explainable AI and question-driven explanations
This work is situated in the field of explainable artificial intelligence (XAI), which aims to make algorithmic models understandable to humans (Ali et al., 2023; Barredo Arrieta et al., 2020; Adadi and Berrada, 2018; Dhanorkar et al., 2021), and, more specifically human-centred XAI (HCXAI), which examines the influence of human factors on user interaction with system explanations (Ehsan et al., 2023). Our study covers two high-risk ”algorithmic decision-making” systems (ADM systems) (European Parliament. Directorate General for Parliamentary Research Services., 2019; Commission, 2021) with humans as decision subjects, i.e., ”affected stakeholders” (Langer et al., 2021).
The term ”stakeholder” describes roles of people involved in the development or deployment of ADM, such as deployers, developers, and affected parties (Langer et al., 2021; Barredo Arrieta et al., 2020). Previous work has established the value of analyzing stakeholders’ information needs (Liao et al., 2020) to use them as basis for the design of explanations (Liao et al., 2021). As explanations of AI systems were originally used by developers to better understand the system’s behavior (Langer et al., 2021) they tend to be targeted towards technical experts, making them less suitable for ”lay people” (Shen et al., 2020; Szymanski et al., 2021; Shulner-Tal et al., 2022) or ”AI novices”. While this work thus follows the concept of question-driven explanations (Liao et al., 2021) and uses a similar approach to Liao et al. (2020)’s collection of AI practitioner’s information needs in the XAI Question Bank, the focus is shifted to another audience. In contrast to more recent work, which investigates how to adapt the XAI Question Bank to end-users (sipos2023), this work aims to provide insight on how explanations can be rendered more accessible and suitable for affected stakeholders, who are not necessarily end-users.
2.2. Perspectives on explanations and understanding
Explanations are a means to increase the recipient’s understanding of a system or decision (Langer et al., 2021), so as to enable actions such as value assessments (Kim et al., 2023), decision-making (Lee et al., 2019), and contestation (Alfrink et al., 2022). As numerous ways exist to define understanding (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Schmude et al., 2023b), we use a definition that aims to preserve the epistemic aspects while allowing for operationalization in an empirical setting. We thus describe understanding as i) connecting and applying information whereas ”knowledge” only stores it (Grimm, 2019; Baumberger et al., 2017), ii) involving both sympathy and causality depending on the object that is understood (Grimm, 2019; Wiggins and McTighe, 2005), and iii) attempting to grasp the underlying structure of a phenomenon (Zagzebski, 2019).
In order to examine how explanations interact with the processes of understanding, we refer to theories by Keil (2006) and Lombrozo and Wilkenfeld (2019), stating that explanations call attention to the incompleteness of one’s understanding, prompting a reaction: (i) filling the gap with further information, (ii) ”outsourcing” it to another mind, or (iii) mentally rendering the gap irrelevant. Developing understanding, in this sense, means filling understanding gaps until one decides to have reached a ”working understanding” (Keil, 2003, 2006) and stops. The ”illusion of explanatory depth” (Rozenblit and Keil, 2002), an overestimation of one’s understanding, can perturb this process and lead to incomplete understanding (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Chromik et al., 2021).
For our analysis, we further use the concept of three ”explanatory stances” (Keil, 2021, 2006; Lombrozo and Gwynne, 2014; Lombrozo and Wilkenfeld, 2019; Dennett, 1998): a theory aiming to capture how people make sense of information and generalize from it. Previous XAI literature used explanatory stances to describe how explanations are framed and processed (Páez, 2019; Zerilli, 2022; Byrne, 2023; Miller, 2019), which we use as basis for our application in an empirical setting. Following, we describe the three explanatory stances in the order mechanical, design, and intentional:
The mechanical stance222Both the terms ’mechanical’ and ’mechanistic’ are used in the literature (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Keil, 2021). Where ’mechanical’ refers to the stance, ’mechanistic’ refers to the explanation. appeals to ”parts, processes, and proximate causal mechanisms” (Lombrozo and Wilkenfeld, 2019) and focuses on low-level technical detail. Explaining why an alarm clock buzzes, for example, in a mechanical stance would state that the power circuit was completed (Lombrozo and Wilkenfeld, 2019). In contrast to functional and intentional explanations, mechanistic explanations do not take into account aims or beliefs (Páez, 2019).
The design stance focuses on functions and goals beyond mechanical interactions in a system (Keil, 2006) and describes why things happen by looking at their aims and purposes (Lombrozo and Gwynne, 2014). Explaining why an alarm clock buzzes in a design stance would state that it functions according to its design: waking its owner at the set time (Lombrozo and Wilkenfeld, 2019; Dennett, 2006).
2.3. Explainability in responsible AI and regulations
Explainability is seen as a pillar of responsible AI (RAI) (Barredo Arrieta et al., 2020; Baeza-Yate and Matthews, 2022; Ehsan et al., 2023), which advocates for fair, accountable, and explainable implementations of AI systems. However, ensuring adherence to RAI principles meets practical challenges. For example, deployments of AI can lead to what Ehsan et al. (2023) and Ackerman (2000) call the sociotechnical gap: ”a divide between the technical affordances and the social needs”. This gap can lead to mismatches between what is expected of a system and what it actually does, and is further aggravated if stakeholders’ interests or values are in conflict (Jakesch et al., 2022).
The draft AI Act, an EU regulation aiming to guide the development and deployment of high-risk AI systems, prescribes measures to uphold RAI principles (Commission, 2021; Panigutti et al., 2023; Sovrano et al., 2022; Veale and Zuiderveen Borgesius, 2021), such as ”the provision of relevant documentation and information to the user” (Panigutti et al., 2023). Which information is considered relevant and how it should be conveyed, however, remains unspecified (Nannini et al., 2023). Lucaj et al. (2023) therefore argue that doubts remain as to how the high-level regulations can improve explainability in practice. To ensure that information is both relevant and intelligible for affected stakeholders, their information needs ought to be elicited and incorporated (Kim et al., 2023). This work aims to contribute to this process by giving an overview of affected stakeholders’ information needs, to be used for explanation design (Madaio et al., 2020; Hagendorff, 2022; Ferrario and Loi, 2022).
3. Method
To elicit information needs of affected stakeholders, we conducted an interview study in German language with 24 participants recruited from a local job agency and a local apartment complex inhabited mainly by retirees (Section 3.4). We presented participants with one of two ADM use cases (Section 3.1): an employment prediction algorithm or a health wristband for elderly care. Participants were tasked to vote whether the presented use cases should be adopted in general. Before deciding, participants were asked to gather information on the systems by asking questions in two 15-minute inquiry phases and received verbal explanations by the main author of this paper in response. The study closed with an interview at around 50 minutes. Interviews were analyzed using a combination of inductive and deductive analysis (as described in Section 3.3).
With this procedure, we aimed to fulfill multiple study goals: firstly, we introduced a task that motivated inquiry of information (i.e., voting on adoption), secondly, by first letting participants inquire freely and flexibly about the use cases and second providing the XAI Question Bank (Liao et al., 2020) as reference we could examine both intuitive information needs as well as the bank’s utility for their inquiry, and third, by specifically eliciting participants’ understanding and perceptions we gave room for articulations about aspects not covered by the inquiry itself. The study setup is described in Section 3.2 and the procedure is depicted in Figure 1.
3.1. Use cases
Choice of use cases. Use cases were selected to adhere to the definition of ”high-risk” ADM systems (Commission, 2021) while differing in several dimensions, including: a) domain of use (employment and health), b) degree of tangibility (the wristband is a physical device, while the AMS algorithm does not have a physical counterpart), c) user groups, and d) functionality. These differences were intended to allow for examination of their influence on participants’ information needs (RQ1), understanding (RQ2), and perceptions of the systems (RQ3). As the two use cases are representative of two types of well-known ADM systems that have already been implemented in comparable forms (Niklas et al., 2015; Scott et al., 2022; Kutsarova, 2020; Sahlen et al., 2018), our study aims to provide insight on how to approach explanation design for the future deployment of similar systems.
Use Case A: AMS algorithm. The AMS algorithm 333AMS stands for the Public Employment Agency (Arbeitsmarktservice). is a scoring system meant to predict the employability of job-seekers in Austria. Developed between 2015 and 2021 by a private company for the Austrian Public Employment Agency and piloted in 2018, it was never used as a live system and was put on hold in 2021 due to privacy objections (Allhutter et al., 2020b). The system was meant to use job-seekers’ personal features such as age, education, and nationality444The full list of features is given in the supplementary material. to compare them with the time it took people with similar features to find employment in the past. This comparison would produce a short-term and long-term employment score for each job-seeker, which in turn would be shown to the employment agency’s personnel consultant as a decision aid for suitable support measures. Personnel consultants would be able to override the system’s scores, but would need to give a reason for doing so (Allhutter et al., 2020b; Holl et al., 2018).
The AMS algorithm has been academically and publicly discussed with respect to the risks and benefits of its deployment, including detailed reports about the system’s potential to reproduce intersectional discrimination (Allhutter et al., 2020b; Allhutter, 2021; Scott et al., 2022; Lopez, 2019). As the implementation of such a system could have large effects on job-seekers and as comparable systems have been developed in other countries (Niklas et al., 2015; Scott et al., 2022), the AMS algorithm presented a suitable use case for our study.
Use Case B: Health wristband for elderly care. The wristband is an adaptation to wearable alarm systems already in use in care organizations (Chang et al., 2023; Kutsarova, 2020; Kaun and Taranu, 2020; Sahlen et al., 2018). While existent alarm systems often require the wearer to push a button to alert care personnel, the health wristband in our study was described to continuously evaluate movement and acceleration data and raise an alarm when it detected that the wearer tripped or fell. The system would thereby not require manual activation by the wearer, making it viable for use in private homes. Further, the wristband collects data on the wearer’s sleeping, movement, and eating behavior, which the care facility could monitor and use to improve the organization of care.
3.2. Study setup
3.2.1. Scenario and decision task
Participants were initially informed that they would vote over the ADM system’s adoption after the inquiry phases, introducing an incentive to gain information. After participants were given a consent form and a questionnaire on demographical information and previous knowledge, participants were presented with the decision scenario, which included a brief description of the use case and its planned deployment. The description rendered the participants as part of a small citizen referendum asked to provide their vote on the system. A small box was printed below the scenario description for participants to fill in ”yes” or ”no” after the inquiry phases, acting like a voting ballot. The design was inspired by ballots used in popular votes in Switzerland (Council, 2023).555Popular votes in Switzerland require the voter to write ”Yes” or ”No” on the ballot, which can be argued to be more deliberate compared to a tick box.
3.2.2. Inquiry phases I & II
After being briefed about the decision scenario, participants had 30 minutes in total to ask questions to the study examiner and receive verbal explanations about the ADM use case at hand before taking the decision. The time in which participants could inquire about the use cases was split into two phases of 15 minutes each: In the first phase, participants inquired about the use case intuitively, having received no initial information beyond the scenario description. In the second phase, participants were provided with a reduced version of the XAI Question Bank (Liao et al., 2020)666The amount of questions compiled in the XAI Question Bank (Liao et al., 2020) would have made its application in the study impractical due to the time necessary to scan and understand all items. Unsuitable questions were thus omitted and similar questions summarized. A detailed tabular explanation for the reduction is given in the supplementary material., a collection of questions on AI systems sourced from AI developers and designers, which they could use as a reference for further questions. As the original XAI Question Bank includes a comprehensive collection of questions, we reduced it by summarizing questions that aim towards similar topics. During the inquiry, participants were provided with a pen and paper to take notes.
3.2.3. Verbal explanations
Participants received verbal explanations from the study examiner as responses to their questions. The explanations were based on information prepared in advance and available to the study examiner during the study. For each use case, two pilot studies were conducted to collect a stock of questions that would likely be asked in the subsequent studies and which served as guidance for preparing information. When questions could not be answered due to a lack of information, they were re-formulated to inquire about how and why the requested information would be relevant to the information needs of participants (for examples of interview excerpts, please refer to the supplementary material).
3.2.4. Self-reported understanding, confidence, and risks & benefits
Participants reported their understanding, decision confidence, and perceptions of the system’s risks and benefits (and an explanation for their answer) at two points: first, before the inquiry phases, and second, after making the decision. Self-reported understanding and decision confidence were elicited through 5-point Likert scales. To encourage participants to speak about the trade-offs between the systems’ risks and benefits, these were split into two dimensions: societal and personal. For each elicitation of risks and benefits, two visual analogue scales (10 cm) (Rugg and Petre, 2007) were combined to form a two-dimensional Cartesian coordinate system with the origin located at 5 cm on both axes. By denoting perceived risk on the x-axis and perceived benefit on the y-axis, the coordinate system formed four quadrants of varying risk and benefit combinations (as depicted in Figure 3). Participants drew marks on the axes that were measured and then rounded to the first decimal place.
3.3. Analysis
For RQ1-Information Needs, we use thematic analysis (Braun and Clarke, 2006) to examine participants’ information needs both inductively and deductively using the XAI Question Bank (Liao et al., 2020). All questions raised by participants in the two inquiry phases were coded according to the category of inquiry. Categories already present in the XAI Question Bank (Liao et al., 2020) were assigned deductively, and topics not yet covered were defined inductively from the data. The XAI Question Bank (Liao et al., 2020) thus served as a starting point for a comparative analysis between participants’ information needs and those of AI practitioners, highlighting similarities and discrepancies between the groups. An intermediary version of the coding was checked by three coders for inter-coder agreement by independent analysis of the transcripts and a subsequent comparison of code categories.
For RQ2-Understanding, we evaluated participants’ responses on the self-reported understanding and decision confidence Likert scales, and used a deductive analysis of explanatory stances and understanding challenges as outlined in Section 2. For RQ3-Perception, we analyze how explanations provided during the inquiry phases impact participants’ perceptions of the systems’ risks and benefits, and how these perceptions in turn influence which questions are asked.
3.4. Participants
Table 1 provides an overview of the study participants. We recruited participants from two main locations:In the employment prediction use case, we collaborated with the local employment agency Job-TransFair, which established contact with both job-seekers and personnel consultants. In the wristband use case, we used street sampling in an apartment complex known for the quality of living it offers to retirees.777While we contacted many institutions active in geriatric care work, none wanted to collaborate. The reasons for this are unknown. However, the difficulty of recruiting retirees for study participation over official channels is noteworthy. The recruitment criteria were i) no or little previous experience with either algorithms or AI systems (reported in the questionnaire) and ii) identification with the ”affected stakeholder” group as described below. All interviews were conducted in person either in office spaces or cafés, except for one conducted online. Participants were compensated with 12€ in cash, which was disclosed only after completion of the study so as not to incentivize participation solely for compensation.
All participants of our study can be defined as affected stakeholders (Langer et al., 2021) to varying degrees. In the employment prediction use case, job-seekers were defined as affected if they were searching for employment at the time of the interview or had been searching in the past. Personnel consultants fulfilled multiple stakeholder roles: first, all personnel consultants also had been job-seekers in the past, and second, being the system’s end users, its deployment would affect their interpersonal relations and collaboration with job-seekers. In the wristband use case, the same categories applied regarding geriatric care. Here, no participants were in care currently or in the past, but all had close relatives in care. Further, the age of participants in this use case was required to be around 60 years or older (B4 and B9 are the exception, as their domain expertise as nurses proved relevant to compare information needs and understanding).
| ID | Age | Education | Occupation | Decision | ID | Age | Education | Occupation | Decision |
| A1 | 54 | University | Personnel consultant | Yes | B1 | 62 | A level | Retired | Yes |
| A2 | 61 | Vocational university | Personnel consultant | No | B2 | 68 | A level | Retired | Yes |
| A3 | 54 | Secondary school | Job-seeking | Yes | B3 | 57 | Secondary school | Retired | Yes |
| A4 | 43 | A level | Personnel consultant | Yes | B4 | 46 | Apprenticeship | Geriatric care | No |
| A5 | 27 | Apprenticeship | Job-seeking | No | B5 | 80 | Secondary school | Retired | Yes |
| A6 | 53 | University | Personnel consultant | No | B6 | 68 | University | Retired | Yes |
| A7 | 63 | University | Personnel consultant | n/a | B7 | 58 | University | Freelancer | Yes |
| A8 | 56 | Apprenticeship | Job-seeking | No | B8 | 70 | Vocational university | Retired | Yes |
| A9 | 50 | A level | Job-seeking | No | B9 | 36 | University | Geriatric care | Yes |
| A10 | 63 | Secondary school | Cleaning service | No | B10 | 66 | Secondary school | Retired | Yes |
| A11 | 40 | A level | Personnel consultant | Yes | B11 | 89 | Apprenticeship | Retired | Yes |
| A12 | 25 | Secondary school | Job-seeking | Yes | B12 | 60 | University | Doctor | Yes |
4. Results
In this section, we present our results, structured according to our research questions: information needs of affected stakeholders (RQ1-Information Needs, Section 4.1), participant understanding and decision confidence (RQ2-Understanding, Section 4.2), and perceived risks and benefits (RQ3-Perceptions, Section 4.3).
| Category | Subcategory | Question | Information provided | ||
| Data |
Type |
What kind of data was the system trained on?* |
Content and structure of training data |
Source |
What is the source of the training data?* |
Collection and storage of training data |
Attributes |
Are the data correct / representative / safe? |
Quality, biases, safety of training data |
| System context |
Intention |
What is the intention of deploying it? |
Design idea and underlying motivation |
How does the deployment process work? |
Process of preparing the system for operation |
Consequences |
What are the consequences after deployment? |
Possible effects of deployment |
Development |
How is the system developed? |
Development process and institution |
Who is the system’s intended target group? |
Scope and features of the target group |
Responsibility |
Who is responsible for the system’s deployment? |
Responsible actors |
Perspective |
What do other people think about the system? |
Votes and considerations of others |
Values |
How are moral values considered in the system? |
Alignment with moral principles |
| System usage |
Tool |
How is the system operated? |
Usage in practice, everyday application |
How will the system impact interpersonal relations? |
Interpersonal work, substitution of people |
How is the system integrated into existing structures? |
Interfaces, processes, organization’s influence |
Misuse |
How can the system be misused? |
Scenarios of misuse |
Example |
How would the system handle [this case]? |
Specific example walkthrough |
Costs |
What are the costs? |
Financing, personal costs |
| System specs. |
Features |
What features does the system consider and why?* |
Information fed into system |
Functionality |
What is the scope of the system’s capability?* |
Supported range of functions |
Output |
What kind of output does the system give?* |
Information received from system |
Learning |
How does the system learn?* |
Automated improvement of system |
Logic |
What is the system’s overall logic?* |
General process overview |
Algorithm |
What kind of algorithm was used?* |
Specific model used |
Mistakes |
What kind of mistakes is the system likely to make?* |
Expected errors |
Limitations |
What are the limitations of the system?* |
Known boundaries |
Reliability |
How reliable are the predictions?* |
Dependability of output |
Reflexive |
What does [a machine learning concept] mean? |
ML concept information |
4.1. RQ1-Information Needs: What are affected stakeholders’ information needs when deciding over the adoption of an ADM system?
To address RQ1, we coded participants’ questions raised in the inquiry phases according to reoccurring categories. An overview and diagram are provided in Figure 2, a tabular representation in Table 2. The categorization covers the categories data, system context, system usage, and system specifications, each containing subcategories with a summarizing question. The subsequent sections (4.2, 4.3) describe how this information influenced participants’ decisions on the system’s adoption. The bracketed numbers denote the total number of questions asked in each subcategory.
4.1.1. Data: Type, source, attributes.
Questions on the data used by the system emerged in both use cases (43 questions total) and occurred mostly in the second phase of inquiry after reference to the XAI Question Bank (Liao et al., 2020). The question bank already contains questions about data type (What kind of data was the system trained on?) and data source (What is the source of the training data?). As multiple participants asked about data privacy and misuse (11 questions by 8 participants), we add a subcategory on data safety: Are the data safe?
4.1.2. System context: Intention, deployment process, consequences, development, target group, responsibility, perspective, and values.
This category captures questions that aimed to elicit background information on various facets influencing the system’s development and deployment (69 total).
Intention: What is the intention of deploying it? Inquiries about the intention of the system’s deployment emerged in both use cases, with a majority in employment prediction (12 of 18 total). These questions mostly aimed at the intention of deploying the system, which we differentiate regarding i) practical intention, such as data collection, ”How did they come up with the idea of GPS? What’s the intention behind it?” (B4); and ii) structural intention, such as societal benefit, ”What is the benefit for society? What did they expect from this algorithm?” (A1).
Deployment process: How does the deployment process work? Inquiries about the general deployment process emerged mostly in the employment prediction use case (13 of 19 total), including how the system should be made available, if it had already been tested, and how comparable systems were deployed in other countries.
Consequences: What are the consequences after deployment? Inquiries about effects that could follow from the system’s deployment emerged in low counts in both use cases (9 total). Topics of inquiry included, e.g., the effect on the available workforce and care ratio, and often had a focus on risks and challenges:
Won’t I lose labor potential if I only decide based on educational background and so on? If this algorithm is introduced, doesn’t it lead to the loss of workforce? (A1)
Development: How is the system developed? Questions on the system’s technical development emerged mainly in the employment prediction use case (8 of 9 total) and covered, e.g., the contracted company, duration of development, and the collaboration between developers and domain experts. These questions were asked primarily by personnel consultants, who related the information to their own experience with contracted software developers and compared it to known procedures.
Target group: Who is the system’s intended target group? Questions on the system’s target groups emerged mainly in the wristband use case, were low in number (8 of 10 total), and asked mostly about the individual background of the target group. In contrast, the 2 inquiries in the employment prediction use case focused on target groups on a societal level: ”Would the system be applied for every population group? Can we say that every citizen would then be covered?” (A11).
Responsibility: Who is responsible for the system’s deployment? Four questions, two in each use case, addressed who would be responsible for the system’s deployment. This low count is noteworthy, especially compared to the number of questions on the employment prediction’s societal and political effects observed in other subcategories.
4.1.3. System usage: As a tool, as a sociotechnical component, as part of an organization, misuse, examples, costs.
Much of participants’ information needs pertained to how the system would be used in practice (89 questions in total by over 20 participants). We differentiate questions on the system’s usage into everyday usage as a tool, usage as a sociotechnical component in human interaction, and usage embedded in a larger organizational structure. We include questions on misuse, costs, and examples, as they also address the system’s concrete application.
Usage as a tool: How is the system operated? Questions on practical details of usage were almost exclusively gathered in the wristband use case (30 of 31 total), with only one question emerging in the employment prediction use case. Questions typically concerned how the wristband would be used by family members, including waterproofing, charging, operation, and similar. Questions like these would likely be covered by a manual accompanying the wristband.
Usage as a sociotechnical component: How will the system impact interpersonal relations? Questions addressing how the systems would impact interpersonal work were gathered mostly from the employment prediction use case (13 of 14 total). Participants’ inquiries focused on how the system would affect interaction between consultant and job-seeker and covered both the consultant’s and job-seeker’s personal experience as well as their relationship:
Some people open up only after the second, third, or fourth conversation. […] Here, it’s really about subjective perception. What do I learn from the interview? What do I learn from the person? And what do I make of it? With the algorithm, I would need to know that this dialogue is still possible. (A2)
Usage as part of an organization: How is the system integrated into existing structures? Questions that focused on the system’s interaction with existing organizational structures were present in both use cases, with a majority in the wristband use case (19 of 28 total). Questions on the institutional embedding covered a wide range of topics, including data processing, data access, information flow, contestability, dependence, and control: ”What are the options for correction? Is there a system loop that handles this?” (A7).
Misuse: How can the system be misused? A noticeably low number of questions (2 total) addressed the potential of misusing the system; both emerged in the employment prediction use case. While one participant asked about using the system to commit welfare fraud, the other was concerned about the system being used by the police.
Examples: How would the system handle [this case]? Questions in this subcategory aimed to examine how the system would behave given a specific example case and emerged in low count in both use cases (6 total). Half of these were formulated as: ”How would the algorithm handle my case?” (A2).
Costs: What are the costs? Questions about costs pertained to both personal costs (”What would it cost me?” (B5)) as well as organizational costs (”If the state pays, what is the cost? Is the money well invested?” (A4)). This subcategory appeared more often in the wristband use case (6 of 8 total).
4.1.4. System specifications: Features, functionality, performance, output, predictions, logic, reflexive inquiry.
This category covers all questions focusing on specific aspects of the ADM system and has the highest question count of all categories (187 total). In contrast to the previous categories, many questions on system information emerged only after participants were provided with the XAI Question Bank (Liao et al., 2020) in the second inquiry phase. We focus on describing questions about the system’s features, functionality, performance, and reflexive inquiry, as the remaining subcategories mostly contain questions adopted verbatim from the XAI Question Bank.
Features: What features does the system consider and why? Questions on the system’s input features emerged almost exclusively in the employment prediction use case (28 of 31 total) and covered a range of different inquiry forms. Most questions on features emerged in the first inquiry phase but mirrored questions covered in the XAI Question Bank (Liao et al., 2020), e.g., Which features are considered?, How is a certain feature weighed?, and Why is this feature included? This points to shared information needs between ADM novices and experts, as questions on features in the question bank were sourced from stakeholders with technical knowledge. However, the use case appears to impact strongly whether features are inquired at all.
Functionality: What is the scope of the system? Can it do…? Questions on the system’s functionality emerged almost exclusively in the wristband use case (30 of 34 total) and in the first inquiry phase. The ’functionality’ subcategory differs from the ’usage as a tool’ subcategory in the inquiry’s focus: Questions on functionality inquired about the general existence or workings of specific functions (such as sleep monitoring or heart frequency measurements), whereas questions on the system’s usage as a tool inquired how the system would be operated in practice.
Performance. Questions in this subcategory belonged to one of three categories which are covered by the XAI Question Bank (Liao et al., 2020): errors (What kinds of mistakes is the system likely to make?), limitations (What are the limitations of the system?), and reliability (How accurate/precise/reliable are the predictions?). While questions on reliability only emerged after reference to the question bank, questions on limitations and errors emerged organically and frequently even in the first inquiry phase: ”I mean, the question is, if the system really does make mistakes, how quickly can you give the feedback: ’Oops, that wasn’t a fall, everything’s okay.’ Is that possible?” (B4).
Reflexive inquiry: What does [a machine learning concept] mean? This subcategory captures inquiries about the meaning of questions included in the XAI Question Bank (Liao et al., 2020), which emerged in both use cases (14 total). Often, these took the form of What does [a machine learning terminology] mean?, prompted by terms such as algorithm, biases, and predictions. However, a lack of understanding in vocabulary often correlated with a lack of understanding of the corresponding concept: ”’What kind of data was the system trained on?’ I don’t understand the ’trained’ part. That’s just an input, isn’t it?” (A3). We thus adapt the question slightly for this subcategory to ask about machine learning concepts.
4.2. RQ2-Understanding: How do question-driven explanations interact with participant understanding?
This section provides an analysis of participant understanding: Section 8 describes the effects of inquiry and verbal explanations on self-reported understanding and decision confidence, and Section 4.2.2 describes how participant understanding interacted with question-driven explanations with a focus on explanatory stances and challenges in understanding.
4.2.1. How inquiry impacts understanding and confidence.
Participants reported their perceived understanding and decision confidence on 5-point Likert scales before and after the inquiry phases.888One participant decided to not vote on the use case and did not answer the second query out of concern about how the study results would be used.. We describe three main aspects of these self-reports: the overall level of understanding, changes in understanding, and interactions between understanding and confidence.
| Employment prediction use case | A6 | A11 | A4 | A5 | A8 | A1 | A3 | A10 | A2 | A12 | A9 | A7 |
| Self-reported understanding | ++ | + | = | = | = | = | = | = | - | - | - - - | n/a |
| Confidence | + | = | + | - - | = | + | = | = | = | = | = | n/a |
| Health wristband use case | B8 | B10 | B2 | B4 | B5 | B6 | B7 | B9 | B12 | B1 | B3 | B11 |
| Self-reported understanding | + | + | = | = | = | = | = | = | = | - | - | - |
| Confidence | - | + | + | = | = | = | ++ | = | - | ++ | + | + |
Self-reported understanding was high overall, but ’high understanding’ differs between participants. 3 of 24 participants indicated ”3” or lower in their first self-reported understanding (3 = indifference between understanding and no understanding, 5 = no understanding at all) and only 1 participant indicated ”3” in their second report. However, when participants explained why they perceived their understanding to be high, the level of detail in their accounts differed noticeably. For example, domain expert A2 explained: ”The system should aid in deciding which measure to offer to a job-seeker. […] The question is how the algorithm is created, but the basic classification, I think, makes sense. (A2).
In contrast, participants with less domain expertise tended to provide less detailed explanations. B5, for example, referred to their reception of AI in the media as a reason for their high self-reported understanding: ”There is a lot of talk about artificial intelligence in the media. Mostly at night. I watch it when I can’t sleep.” (B5). Other participants spoke about the general employment market (A5), consequences of digitization (A9), and comparable systems (e.g., B8, A12). This illustrates a discrepancy in what people see as ”high” understanding, possibly resulting from the difficulty of assessing one’s understanding (Keil, 2006) (Section 5).
Inquiry phases mostly did not influence self-reported understanding. After proceeding through the inquiry phases and the decision, most participants reported their understanding to have remained the same, 4 participants reported increased understanding and 6 reported declines (as seen in Table 3). As expected, increased self-reported understanding was explained through information gained in the inquiry phases. In contrast, decreases in understanding seemed to have various reasons, e.g., for A2 the decrease in understanding was accompanied by heightened scrutiny:
Interesting, I was so convinced about it at first, but now there’s something missing from my point of view. Basically, I’m still convinced that it’s a good thing, but it’s not ready yet. (A2)
Previous studies have observed that information gain can paradoxically result in decreased perceived understanding, even though objective understanding increased, likely resulting from a new awareness for unknown information (Cheng et al., 2019).
Confidence increased for 9 participants but seems not to depend on understanding. In contrast to understanding, inquiry led to more increases than decreases in perceived decision confidence (as seen in Table 3). A final increase in confidence for several participants seemed to depend on how the system would work in practice: ”By and large, I know how the system works. Otherwise, to be sure, I would have to see how it actually works in practice.” (A4). Participants however offered little explanation as to why understanding did not increase together with their confidence.
4.2.2. How question-driven explanations interacted with participant understanding.
The cognitive sciences outline a number of processes governing how people seek and process information (described in Section 2). In this section, we analyze two main aspects of participant understanding: the role of explanatory stances and the challenges that prevented participants from effectively filling their gaps in understanding.
Explanatory stances guide how information is acquired. To analyze if participants preferred a specific way of seeking information, we apply the concept of the three explanatory stances (mechanical, design, and intentional) (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Dennett, 1998).999A more detailed introduction is provided in section 2. Which stance is chosen guides how information is inquired and understood, e.g., inquiring information from a mechanical stance would prompt explanations about technical details. Examining which stance is applied in inquiry thus indicates if participants often used a particular stance for a particular topic. In the following, we describe which stances were applied most often in the question categories outlined in Section 4.1.
Questions about data and system specifications often applied the mechanical stance. The mechanical stance focuses on parts, processes, and mechanisms and does not consider goals or intentions (Lombrozo and Wilkenfeld, 2019). Participants often applied the mechanical stance when asking about, e.g., features, predictions, and output. This likely resulted from reference to the XAI Question Bank (Liao et al., 2020), which mostly contains questions that elicit mechanistic and functional explanations, such as: ”What is the sample size of the training data?” (A4), ”How does the algorithm actually work?” (A3), and ”How often does the system make mistakes?” (B1). Small differences in question phrasing, however, can lead to different explanations. Compare: ”What is the sample size of the training data?” (A4) and ”The sample size of the data, I assume, is large enough?” (A5). While the first one is taken verbatim from the XAI Question Bank (Liao et al., 2020), the second raises the point: ”Large enough for what?”, which rather applies the design stance in asking for functional information.
Questions about the system’s context and usage often applied the design stance. The design stance focuses on how functions and practical goals are realized in terms of design (Keil, 2006). Participants tended to apply this stance to inquire about, e.g., practical intention, consequences, and usage as a tool or in an organization. Typical design inquiries included: ”What’s the benefit of knowing whether the patient was at the refrigerator at 3:00 or in the bathroom at 2:00?” (B4), and ”What if someone is lying in bed and now has a heart attack? Can the device detect that, too?” (B5). Notably, questions on functionality almost exclusively emerged in the wristband use case.
Questions about discrepancies in values applied the intentional stance. The intentional stance focuses on underlying purpose and connects explanations to the beliefs and desires involved (Keil, 2006; Dennett, 1998). In both use cases, participants applied this stance when asking about the structural intentions of the deploying institutions: ”Would that mean that the employees of the job agency would need to perform better?” (A4). Participants also applied this stance when inquiring about their rights as affected stakeholders: ”If I’m analyzed, I would like to have the results of the analysis. Is there such a thing as a right of appeal?” (A8). Combined, this indicates that the intentional stance was applied when participants thought their own values would not be considered in the system, which seemed to occur more often in the employment prediction than the wristband use case.
Challenges in acquiring information to develop understanding. Explanations bring the incompleteness of one’s understanding into focus and prompt specific reactions: filling the gap with further information, ”outsourcing” it to another mind, or mentally rendering it irrelevant (Keil, 2006). The explanation setting used in this study explicitly supported filling gaps with information by letting participants flexibly inquire about any topic. However, participants also met challenges in acquiring information and developing understanding. In the following, we describe three of these challenges:
Blind spots inhibit further inquiry. In the closing interviews, multiple participants expressed that they did not understand all relevant aspects of the system but could not articulate which information was missing. Reasons given for this effect included a lack of mathematical knowledge (A6), too little preparation (A8), pressure of the study situation (B5, A12), and uncertainty about the system’s practical behavior (A1, A4, B1, B9): ”Experience has taught me that it’s always the practical application that raises issues.” (B1). We call understanding gaps that could not be closed ”blind spots”.
Outsourcing understanding divides cognitive labor but risks over-dependence. Multiple participants asked about other people’s judgment of the system, e.g., the job center’s employees (A8), digital natives (A9), their families (B11), and the study examiner (A11). This ”outsourcing” (Keil, 2003) is a common process to divide the cognitive labor of understanding between trusted people. Critically, some participants stated intention to outsource their whole understanding: ”The first question is: Would you have said yes to this system?” (A9). Trust is thus a crucial part of stakeholders’ understanding when outsourcing is used, especially as it risks over-dependence on the entrusted person.
Abandoning of understanding undermines the explanation effort. Some participants expressed that they would likely not care about the system even if it was used on them, resulting from a lack of interest or the belief that their voice would be irrelevant: ”I still think that it’s not my interest that is being represented here, but actually the interest of the job agency.” (A8). A pronounced disinterest in topics despite their relevance for oneself is known in political science as ”civic ignorance” (Lupia, 2016)101010The term describes that ignorance on public matters applies to almost everyone (Lupia, 2016).. Research has been conducted on ways to remedy this disinterest and will be briefly discussed in Section 5.
4.3. RQ3-Perception: How do question-driven explanations impact perceptions of the systems’ risks and benefits?
Participants were asked for their perceptions of risks and benefits before inquiry and after making the decision (Section 3). Changes in their perceptions from the first to second report are depicted in Figure 3, showing strong changes in the employment prediction use case and less so in the health wristband use case. In the following, we provide a brief overview of participants’ perceptions in both use cases.
Perceptions changed strongly in employment prediction use case. Participants in the employment prediction use case showed diverse initial perceptions (as seen in Figure 3) and strong changes in their perceptions after the inquiry phases. Perceptions were further likely influenced by the participants’ familiarity (Kramer et al., 2018) with the AMS algorithm, which drew considerable public attention during its piloting phase in 2018 (Allhutter et al., 2020a). Frequently mentioned benefits included a potentially reduced workload, debureaucratization, and guidance for personnel consultants, while frequently mentioned risks included a lack of care for the individual, fear of discrimination, and overreliance on algorithmic decisions. A6, for example, reported a decrease in perceived societal risk and benefit, while also reporting an increase in perceived personal benefit, explaining:
My intuition is that an algorithm like this can make my work a lot easier because it gives me guidance. But if it’s used too much as a guideline, then it naturally restricts the options of the people I look after. (A6)
In contrast, A12 reported a strong decrease in perceived risk and increased benefit in both dimensions, stating:
Honestly, I haven’t heard anything bad. It can only have positive benefits. […] I say it’s worth a try. And in any case, an assessment will be made based on the data that is available. (A12)
In total, we observe mixed effects that question-driven explanations had on participants perceptions in the employment prediction use case. While the provided information certainly enabled a shift in perspective, the trajectory of that shift seems to differ heavily, even though the information provided did not change between participants except, of course, if different questions were asked.
Perceptions changed less in the health wristband use case. Participants in the health wristband use case reported less pronounced changes in perceptions and overall saw the use case rather clearly as one of low risk and high benefit. Often participants weighed the risk of surveillance against the benefit of the wearer’s safety, as B11 explains:
For me I don’t see any disadvantages. But I could imagine that there are people who say: ”I don’t want to be monitored. Even if I fall down and nobody comes and I bleed to death, then I’ll just bleed to death.” (B11)
B12 assigned the system a higher societal risk but also a high benefit, concluding:
The risk of data collection is a risk inherent to the system. Either you accept this risk and have the benefit of these systems, or you do without it. You probably can’t have one without the other. (B12)
5. Discussion
In this section, we discuss the findings of our study regarding information needs and understanding (Section 5.1), stakeholder aims and explanatory stances (Section 5.2), and measures to incorporate affected stakeholders into explanation design (Section 5.3), and provide a list summarizing the implications of our findings (Section 5.4).
5.1. Eliciting affected stakeholders’ information needs and resolving understanding issues
One of XAI’s enduring key questions is: What makes a ”good” explanation? Recent work outlines that, to produce understanding, explanations should be contrastive, selected, and social, while being first and foremost contextual (Miller, 2019), that they should interface directly with people’s causal reasoning (Byrne, 2023) and consider the learner’s aim and prior knowledge (Mueller et al., 2019). These demands can be realized practically by identifying relevant information needs and examining how this information is processed. In the following, we present how the findings of our study support this analysis.
5.1.1. Collecting information needs and using them as reference for explanation design.
The findings described in Section 4.1 show that affected stakeholders’ information needs cover all parts of an ADM system’s ”lifecycle” (Dhanorkar et al., 2021), ranging from technical details over deployment consequences to different dimensions of usage. Our proposed ”XAI Novice Question Bank” (Figure 2) aims to provide a summary of these topics which can be used for future explanation approaches. In empirical settings, presenting participants with collections like the XAI Novice Question Bank and the XAI Question Bank (Liao et al., 2020) can support the ”intelligibility” (Lim and Dey, 2009) of a system and reduce the risks of ”blind spots” , i.e., information that stakeholders cannot designate or locate.
To ascertain which topics are actually deemed relevant in a given use case, however, requires speaking to the people involved. This is demonstrated by the differences in information needs between use cases A and B as well as the differences to Liao et al. (2020)’s study. In line with previous work, we observe that various contextual factors mediate which information need participants articulate, including: familiarity with the presented system and comparisons with similar systems (Kramer et al., 2018), tangibility of the system (Long et al., 2021), perceived risks and benefits (Araujo et al., 2020) (e.g., trade-off between safety and privacy in the wristband use case), domain expertise (Wang and Yin, 2021) (e.g., questions about development), and technical expertise (Cheng et al., 2019).
5.1.2. Resolving understanding challenges in question-driven explanations.
Our findings highlight two key issues of how on-explanations and understanding interact: First, letting stakeholders acquire information independently supports personalized explanations (Conati et al., 2021), but at the same time transfers to stakeholders the burden of deciding when a sufficient understanding has been reached, which can be difficult to assess (Keil, 2006). In consequence, self-reports of understanding can be unreliable and prone to the ”illusion of explanatory depth” (Rozenblit and Keil, 2002; Chromik et al., 2021): an overestimation of one’s understanding. Previous work has outlined various strategies to offset these issues, including deliberate self-explanation (Duckworth, 2001), better initial introductions to the system (Chromik et al., 2021), and actively checking for people’s understanding by letting them predict system outcomes before providing explanations (Buçinca et al., 2020). We argue that further research is needed to determine suitable formats for self-assessments of explanatory understanding in XAI.
Second, several issues can interfere with information acquisition, including blind spots and outsourcing. We use the term blind spots for information that participants missed but could not articulate, which might correspond to a lack of ”intelligibility types” (Lim and Dey, 2009), i.e., a lack of knowledge about available information and ways to inquire it. References like our proposed XAI Novice Question Bank (Figure 2) and the XAI Question Bank (Liao et al., 2020) should help in addressing this issue by outlining the relevant categories of information. Outsourcing, on the other hand, relies on entrusting other people to share the burden of understanding, a topic that seems underexplored in current XAI research considering its importance in understanding. We argue that future work should examine group settings and verbal explanation modalities for possibilities of including outsourcing as an understanding strategy.
5.2. Using explanatory stances to address different aims of affected stakeholders
In this section, we discuss how future explanation approaches can incorporate the stakeholders’ aims and explanation’s purpose in the design process by using the concept of explanatory stances (Keil, 2021, 2006; Lombrozo and Wilkenfeld, 2019). We first discuss how participants’ aims emerged in our study and second outline how explanatory stances can be used to convey information according to these aims.
5.2.1. Considering the stakeholder’s aim and the explanation’s purpose.
Explanations have the purpose of enabling stakeholders to attain their aims (Freiesleben and König, 2023), such as calibrating their acceptance or trust towards a system (Shin, 2021), or performing an action, like contesting a decision or requesting human intervention (Alfrink et al., 2022). Participants in our study were asked to vote over the adoption of the ADM systems to create a purpose which would incentivize acquisition of information through question-driven explanations. This worked for participants who used the explanations to underpin their decision and in turn reported a slight increase in decision confidence (Table 3).
On the other hand, the perceived risks and benefits of the ADM systems seems to have influenced how participants defined their aims and approached inquiry: Participants inquired more about societal and structural factors in the employment use case, and more about practical and operational factors in the wristband use case. Explanation approaches for affected stakeholders should thus consider that the domain of the use case and stakeholders’ perceptions towards it impact which information is deemed relevant. Acknowledging the large role that the institutional context plays in how affected stakeholders perceive ADM (Brown et al., 2019; Woodruff et al., 2018; Lee et al., 2019; de Fine Licht and de Fine Licht, 2020; Veale et al., 2018), explanations likely cannot always meet the aims of all stakeholder groups, especially when conflicts of interest exist.
5.2.2. Explanatory stances can be useful to guide information acquisition.
People can differ considerably in how they approach information acquisition and understanding, depending in part on their stakeholder role and their aims. As explanatory stances (mechanical, design, and intentional (Dennett, 1998)) describe the ways in which people make sense of information and generalize from it, applying them in explanation design could help to convey information more effectively. Previous XAI research considered explanatory stances in conceptual approaches (Páez, 2019; Miller, 2019; Zerilli, 2022), but rarely empirically. We argue that future work in XAI should investigate explanatory stances for two main reasons:
First, our interviews showed that certain topics lend themselves for the application of certain stances. For example, participant inquiries about system specifications mostly applied a mechanical or design stance, while inquiries about system context rather applied the design and intentional stance (Section 4.2.2). Which stance is applied can depend on the stakeholder role, for example: Features of an ADM system can be relevant for both developers, who could inquire about the features’ data types, and affected stakeholders, who could inquire how features capture a real world context. Importantly, both stakeholders could ask the same question: What features does the system consider and why? Considering which stance a stakeholder likely applies in acquiring information might therefore allow for better personalization of question-driven explanations.
Second, encouraging application of different explanatory stances can support a change in perspective. Work in the cognitive sciences describes the possibility that the design stance might be preferred over the mechanical stance by default (Lombrozo and Wilkenfeld, 2019).111111Preference in explanatory stances is a controversial discussion, the literature currently does not draw general conclusions (Keil, 2021; Lombrozo and Wilkenfeld, 2019; Kelemen et al., 2013). In our study, we observe that both the design and intentional stance in fact emerged more frequently than the mechanical stance in the first inquiry phase. Therefore, explanation approaches can encourage an active change in perspective by providing material that uses a different stance than the participant and steers attention to information that is not explored autonomously. For example, introducing the XAI Question Bank (Liao et al., 2020) in the second inquiry phase of our study incited participants to apply a mechanical stance, which they did less in the first (unguided) phase of inquiry.
5.3. Taking steps to incorporate affected stakeholders into human-centered XAI
Affected stakeholders, as decision subjects, are arguably the most vulnerable stakeholder group of ADM and should therefore be able to receive explanations about systems that might affect them (European Parliament. Directorate General for Parliamentary Research Services., 2019; Commission, 2021; Vredenburgh, 2022; Panigutti et al., 2023). Our findings draw attention to two challenges in using question-driven explanations for this purpose: First, participants in our study expressed interest to learn about intentional information, which is not typically covered in explanations. Second, affected stakeholders might have no interest in interacting with explanations or in taking part in decisions about ADM. In the following, we briefly discuss these issues with a perspective on future work in XAI research.
5.3.1. Handling explanations of intention.
In our study, many participants applied the ”intentional stance” (Dennett, 1998) to inquire about the ADM systems’ context and usage, aiming to examine the beliefs and values ingrained in the systems (Section 4.2). These inquiries were often motivated by how participants perceived the deploying institution and the involved risks and benefits (Brown et al., 2019; Woodruff et al., 2018). However, explaining the intentions and values of a system’s deployment means taking a stance towards the values involved. Further, intentional information is not strictly factual and thus might not be easily accessible. Then again, human explanations are by nature social (Miller, 2019) and transactional (Keil, 2006), and tend to use subjective views to better convey information.
Consequently, determining the explanation’s purpose and stakeholders’ aim might facilitate the decision of whether intentional information should be included (Freiesleben and König, 2023). For example, if the explanation’s purpose is a solely mechanical or functional description of the system in order to enable operation, intentional information is likely negligible. In contrast, when stakeholders should form value assessments or investigate possibilities for contestation (Alfrink et al., 2022), intentional explanations can contribute meaningful information, e.g., by using ”justifications” (Biran and Cotton, 2017; Biran and McKeown, 2017; de Fine Licht and de Fine Licht, 2020) to explain on which grounds decisions are made. As stakeholders’ interest in the beliefs and values of ADM is a central information need, we argue that future research should examine how this information can be conveyed meaningfully.
5.3.2. Taking steps to connect explainability to civic education.
Our studies show that affected stakeholders are not always interested in understanding ADM, even if it might affect them. Assuming that one goal of explainable AI is to provide the most information to the most people in order to increase understanding of public AI systems, the question of how to redeem the interest of these stakeholders is central. Work on similar issues in democratic participation terms this issue ”civic ignorance”: people’s general disinterest in laws and regulations that shape their lives (Lau and Redlawsk, 2001; Lupia, 2016). In fact, civic ignorance is not meant as a stigma, but as a common effect that applies to most people, including political experts (Lupia, 2016). The corresponding research describes viable remedies, beginning with identifying the intersection of i) information that the public is interested in because it advances their core concerns and ii) information that is sufficient to enable performance of a task, such as judging a political topic. When this intersection, the ”information that matters” (Lupia, 2016), has been outlined, subsequent steps can develop educational measures that have better chances to redeem public interest.
Regarding XAI, this paper aims to accomplish the first step of identifying information that is both relevant to affected stakeholders and sufficient to form a judgment. Recent work in AI literacy has developed catalogs of competences and design considerations for educational interventions on AI (Long and Magerko, 2020; Ng et al., 2021). We argue that future work should combine these empirical and conceptual insights to design explanations that effectively convey information to affected stakeholders.
5.4. Implications
In this section, we briefly summarize the implications that our findings and discussion have for the design of explanations for affected stakeholders.
Affected stakeholders’ information needs are diverse and differ between use cases. Our proposed XAI Novice Question Bank (Figure 2) extends the well-known XAI Question Bank (Liao et al., 2020) by including questions on the system context and usage. It can thus be leveraged to design better-targeted explanations for affected stakeholders.
Affected stakeholders’ information needs depend on the actions they want to take, such as assessing values, voting, or contestation. Explanations should support affected stakeholders in realizing their specific aim(s).
Question-driven explanations allow stakeholders to decide when their understanding is sufficient for their aims. But assessing one’s own understanding is inherently difficult and prone to cognitive biases (Section 5.2). Facilitating this assessment could be necessary to avoid the illusion of explanatory depth in question-driven explanations.
Information can be understood from different angles. Certain angles might be more intuitive for certain stakeholders: domain experts might ask about a feature’s purpose (functional) while job-seekers might ask why it was included (intentional). The chosen angle can impact which explanation is suitable.
Information on the underlying reasoning of why an ADM system is deployed can be relevant for affected stakeholders. Typical explanations focus on the workings of the system rather than this type of contextual information. XAI research should also provide guidelines on how the decision-making behind a system’s development and deployment can be explained.
Researchers developing explanations for affected stakeholders ought to be aware that people might be disinterested in explanations despite being affected, caused, for example, by a perceived inability to understand ADM or a feeling of powerlessness. Future research should examine further how these factors impact explanations for affected stakeholders.
6. Limitations
Our participants were recruited from a local job agency and an apartment complex from the same geographical region, perhaps resulting in regional or cultural biases. Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results. While this limits the generalizability of our results regarding sex and/or gender aspects, it does not limit the overall validity of our results. Further, the articulated information needs and observed processes of understanding are tied to the presented use cases, as demonstrated in Section 4.2, and to the verbal exchange with the study examiner, meaning that participants would likely ask different questions if information was presented in another fashion or by another person. Concerning the size of the participant sample, we are guided by research on qualitative methods, which suggests that 16 to 24 interviews are needed to reach adequate code and meaning saturation (Hennink et al., 2017).
7. Conclusion
The use of public AI systems in society can have considerable impact on people affected by algorithmic decisions. These affected stakeholders should be able to understand AI systems in order to take action, such as contesting decisions or asking for human intervention. Explanations can support this empowerment, but affected stakeholders’ information needs have rarely been covered in explanation approaches. To address this gap, we provide a collection of affected stakeholders’ information needs titled the ”XAI Novice Question Bank” (Figure 2), compiled from interviews with 24 participants and covering two ADM use cases (employment prediction and a health wristband). Our approach aims to extend the concept of question-driven explanations, inspired by the XAI Question Bank (Liao et al., 2020), to lay audiences and highlights the relevance of information needs about an ADM system’s context and usage. Further, we analyze participant understanding by applying cognitive theories to examine from which perspective participants approach information acquisition. We provide suggestions on matching explanations to the way stakeholders make sense of information and on using question-driven explanations to overcome understanding challenges. Lastly, we examine participants’ perceptions of the systems’ risk and benefits, finding that higher perceived risk likely prompts participants to inquire about intentional information, such as the motivation and consequences of deploying a system. We close with six key implications that our findings have for the design of explanations for affected stakeholders and posit that more research is needed to create suitable explanations for this crucial stakeholder group.
Acknowledgements.
This work has been funded by the Vienna Science and Technology Fund (WWTF) [10.47379/ICT20058] as well as [10.47379/ICT20065]. We further thank the Vienna Job-TransFair team for their valuable contributions.References
- (1)
- Ackerman (2000) Mark S. Ackerman. 2000. The Intellectual Challenge of CSCW: The Gap Between Social Requirements and Technical Feasibility. Human–Computer Interaction 15, 2-3 (2000), 179–203. https://doi.org/10.1207/S15327051HCI1523_5
- Adadi and Berrada (2018) Amina Adadi and Mohammed Berrada. 2018. Peeking Inside the Black-Box: A Survey on Explainable Artificial Intelligence (XAI). IEEE Access 6 (2018), 52138–52160. https://doi.org/10.1109/ACCESS.2018.2870052
- Alfrink et al. (2022) Kars Alfrink, Ianus Keller, Gerd Kortuem, and Neelke Doorn. 2022. Contestable AI by Design: Towards a Framework. Minds and Machines (13 Aug 2022). https://doi.org/10.1007/s11023-022-09611-z
- Ali et al. (2023) Sajid Ali, Tamer Abuhmed, Shaker El-Sappagh, Khan Muhammad, Jose M. Alonso-Moral, Roberto Confalonieri, Riccardo Guidotti, Javier Del Ser, Natalia Díaz-Rodríguez, and Francisco Herrera. 2023. Explainable Artificial Intelligence (XAI): What we know and what is left to attain Trustworthy Artificial Intelligence. Information Fusion 99 (Nov. 2023), 101805. https://doi.org/10.1016/j.inffus.2023.101805
- Allhutter (2021) Doris Allhutter. 2021. Ein Algorithmus zur effizienten Förderung der Chancen auf dem Arbeitsmarkt? WISO – Zeitschrift für Sozial- und Wirtschaftswissenschaften 44.JG (2021), 82–95.
- Allhutter et al. (2020a) Doris Allhutter, Florian Cech, Fabian Fischer, Gabriel Grill, and Astrid Mager. 2020a. Algorithmic Profiling of Job Seekers in Austria: How Austerity Politics Are Made Effective. Frontiers in Big Data 3 (2020). https://doi.org/10.3389/fdata.2020.00005
- Allhutter et al. (2020b) Doris Allhutter, Astrid Mager, Florian Cech, Fabian Fischer, and Gabriel Grill. 2020b. DER AMS-ALGORITHMUS: Eine Soziotechnische Analyse des Arbeitsmarktchancen-Assistenz-Systems (AMAS). Technical Report. https://epub.oeaw.ac.at/0xc1aa5576_0x003bfdf3.pdf
- Ananny and Crawford (2018) Mike Ananny and Kate Crawford. 2018. Seeing without knowing: Limitations of the transparency ideal and its application to algorithmic accountability. New Media & Society 20, 3 (March 2018), 973–989. https://doi.org/10.1177/1461444816676645
- Araujo et al. (2020) Theo Araujo, Natali Helberger, Sanne Kruikemeier, and Claes H. de Vreese. 2020. In AI we trust? Perceptions about automated decision-making by artificial intelligence. AI & SOCIETY 35, 3 (01 Sep 2020), 611–623. https://doi.org/10.1007/s00146-019-00931-w
- Baeza-Yate and Matthews (2022) Ricardo Baeza-Yate and Jeanna Matthews. 2022. Statement on Principles for Responsible Algorithmic Systems. https://www.acm.org/binaries/content/assets/public-policy/final-joint-ai-statement-update.pdf
- Barredo Arrieta et al. (2020) Alejandro Barredo Arrieta, Natalia Díaz-Rodríguez, Javier Del Ser, Adrien Bennetot, Siham Tabik, Alberto Barbado, Salvador Garcia, Sergio Gil-Lopez, Daniel Molina, Richard Benjamins, Raja Chatila, and Francisco Herrera. 2020. Explainable Artificial Intelligence (XAI): Concepts, taxonomies, opportunities and challenges toward responsible AI. Information Fusion 58 (2020), 82–115. https://doi.org/10.1016/j.inffus.2019.12.012
- Baumberger et al. (2017) Christoph Baumberger, Claus Beisbart, and Georg Brun. 2017. What is Understanding? An Overview of Recent Debates in Epistemology and Philosophy of Science. In Explaining Understanding: New Perspectives from Epistemolgy and Philosophy of Science, Stephen Grimm, Christoph Baumberger, and Sabine Ammon (Eds.). Routledge, 1–34.
- Biran and Cotton (2017) Or Biran and Courtenay V. Cotton. 2017. Explanation and Justification in Machine Learning : A Survey Or.
- Biran and McKeown (2017) Or Biran and Kathleen McKeown. 2017. Human-Centric Justification of Machine Learning Predictions. In Proceedings of the Twenty-Sixth International Joint Conference on Artificial Intelligence, IJCAI-17. 1461–1467. https://doi.org/10.24963/ijcai.2017/202
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (Jan. 2006), 77–101. https://doi.org/10.1191/1478088706qp063oa
- Brown et al. (2019) Anna Brown, Alexandra Chouldechova, Emily Putnam-Hornstein, Andrew Tobin, and Rhema Vaithianathan. 2019. Toward Algorithmic Accountability in Public Services: A Qualitative Study of Affected Community Perspectives on Algorithmic Decision-making in Child Welfare Services. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300271
- Burns et al. (2023) Alyxander Burns, Christiana Lee, Ria Chawla, Evan Peck, and Narges Mahyar. 2023. Who Do We Mean When We Talk About Visualization Novices?. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 819, 16 pages. https://doi.org/10.1145/3544548.3581524
- Buçinca et al. (2020) Zana Buçinca, Phoebe Lin, Krzysztof Z. Gajos, and Elena L. Glassman. 2020. Proxy tasks and subjective measures can be misleading in evaluating explainable AI systems. In Proceedings of the 25th International Conference on Intelligent User Interfaces. ACM, Cagliari Italy, 454–464. https://doi.org/10.1145/3377325.3377498
- Byrne (2023) Ruth M.J. Byrne. 2023. Good Explanations in Explainable Artificial Intelligence (XAI): Evidence from Human Explanatory Reasoning. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence. International Joint Conferences on Artificial Intelligence Organization, Macau, SAR China, 6536–6544. https://doi.org/10.24963/ijcai.2023/733
- Capel and Brereton (2023) Tara Capel and Margot Brereton. 2023. What is Human-Centered about Human-Centered AI? A Map of the Research Landscape. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–23. https://doi.org/10.1145/3544548.3580959
- Chang et al. (2023) Fangyuan Chang, Britt Östlund, and Sanna Kuoppamäki. 2023. Domesticating Social Alarm Systems in Nursing Homes: Qualitative Study of Differences in the Perspectives of Assistant Nurses. J Med Internet Res 25 (5 May 2023), e44692. https://doi.org/10.2196/44692
- Cheng et al. (2019) Hao-Fei Cheng, Ruotong Wang, Zheng Zhang, Fiona O’Connell, Terrance Gray, F. Maxwell Harper, and Haiyi Zhu. 2019. Explaining Decision-Making Algorithms through UI: Strategies to Help Non-Expert Stakeholders. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300789
- Chromik et al. (2021) Michael Chromik, Malin Eiband, Felicitas Buchner, Adrian Krüger, and Andreas Butz. 2021. I Think I Get Your Point, AI! The Illusion of Explanatory Depth in Explainable AI. In 26th International Conference on Intelligent User Interfaces. ACM, College Station TX USA, 307–317. https://doi.org/10.1145/3397481.3450644
- Commission (2021) European Commission. 2021. Laying Down Harmonised Rules on Artificial Intelligence. https://digital-strategy.ec.europa.eu/en/library/proposal-regulation-laying-down-harmonised-rules-artificial-intelligence
- Conati et al. (2021) Cristina Conati, Oswald Barral, Vanessa Putnam, and Lea Rieger. 2021. Toward personalized XAI: A case study in intelligent tutoring systems. Artificial Intelligence 298 (Sept. 2021), 103503. https://doi.org/10.1016/j.artint.2021.103503
- Council (2023) Swiss Federal Council. 2023. Popular vote. https://www.admin.ch/gov/en/start/documentation/votes.html. Official website of the Swiss Federal Council describing the popular vote. Last accessed: 14.10.2023.
- de Fine Licht and de Fine Licht (2020) Karl de Fine Licht and Jenny de Fine Licht. 2020. Artificial Intelligence, Transparency, and Public Decision-Making: Why Explanations Are Key When Trying to Produce Perceived Legitimacy. AI Soc. 35, 4 (dec 2020), 917–926. https://doi.org/10.1007/s00146-020-00960-w
- Dennett (2006) Daniel Dennett. 2006. Intentional Systems Theory. In The Oxford Handbook of Philosophy of Mind, Brian McLaughlin, Ansgar Beckermann, and Sven Walter (Eds.). Oxford University Press.
- Dennett (1998) Daniel Clement Dennett. 1998. The intentional stance (7. printing ed.). MIT Press, Cambridge, Mass.
- Dhanorkar et al. (2021) Shipi Dhanorkar, Christine T. Wolf, Kun Qian, Anbang Xu, Lucian Popa, and Yunyao Li. 2021. Who needs to know what, when?: Broadening the Explainable AI (XAI) Design Space by Looking at Explanations Across the AI Lifecycle. In Designing Interactive Systems Conference 2021. ACM, Virtual Event USA, 1591–1602. https://doi.org/10.1145/3461778.3462131
- Duckworth (2001) Eleanor Duckworth (Ed.). 2001. ”Tell me more”: listening to learners explain. Teachers College Press, New York.
- Ehsan et al. (2021) Upol Ehsan, Q. Vera Liao, Michael Muller, Mark O. Riedl, and Justin D. Weisz. 2021. Expanding Explainability: Towards Social Transparency in AI systems. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–19. https://doi.org/10.1145/3411764.3445188
- Ehsan et al. (2023) Upol Ehsan, Koustuv Saha, Munmun De Choudhury, and Mark O. Riedl. 2023. Charting the Sociotechnical Gap in Explainable AI: A Framework to Address the Gap in XAI. Proc. ACM Hum.-Comput. Interact. 7, CSCW1, Article 34 (apr 2023), 32 pages. https://doi.org/10.1145/3579467
- Ehsan et al. (2022) Upol Ehsan, Philipp Wintersberger, Q. Vera Liao, Elizabeth Anne Watkins, Carina Manger, Hal Daumé Iii, Andreas Riener, and Mark O Riedl. 2022. Human-Centered Explainable AI (HCXAI): Beyond Opening the Black-Box of AI. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. ACM, New Orleans LA USA, 1–7. https://doi.org/10.1145/3491101.3503727
- European Parliament. Directorate General for Parliamentary Research Services. (2019) European Parliament. Directorate General for Parliamentary Research Services. 2019. Understanding algorithmic decision-making: opportunities and challenges. Publications Office, LU. https://data.europa.eu/doi/10.2861/536131
- Ferrario and Loi (2022) Andrea Ferrario and Michele Loi. 2022. How Explainability Contributes to Trust in AI. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1457–1466. https://doi.org/10.1145/3531146.3533202
- Floridi (2019) Luciano Floridi. 2019. Establishing the rules for building trustworthy AI. Nature Machine Intelligence 1, 6 (June 2019), 261–262. https://doi.org/10.1038/s42256-019-0055-y
- Freiesleben and König (2023) Timo Freiesleben and Gunnar König. 2023. Dear XAI Community, We Need to Talk!. In Explainable Artificial Intelligence, Luca Longo (Ed.). Springer Nature Switzerland, Cham, 48–65.
- Golpayegani et al. (2023) Delaram Golpayegani, Harshvardhan J. Pandit, and Dave Lewis. 2023. To Be High-Risk, or Not To Be—Semantic Specifications and Implications of the AI Act’s High-Risk AI Applications and Harmonised Standards. In 2023 ACM Conference on Fairness, Accountability, and Transparency. ACM, Chicago IL USA, 905–915. https://doi.org/10.1145/3593013.3594050
- Grimm (2019) Stephen R. Grimm. 2019. Varieties of Understanding. In Varieties of Understanding. Oxford University Press, 1–14. https://doi.org/10.1093/oso/9780190860974.003.0001
- Hagendorff (2022) Thilo Hagendorff. 2022. Blind spots in AI ethics. AI and Ethics 2, 4 (Nov. 2022), 851–867. https://doi.org/10.1007/s43681-021-00122-8
- Henman (2020) Paul Henman. 2020. Improving public services using artificial intelligence: possibilities, pitfalls, governance. Asia Pacific Journal of Public Administration 42, 4 (Oct. 2020), 209–221. https://doi.org/10.1080/23276665.2020.1816188
- Hennink et al. (2017) Monique M. Hennink, Bonnie N. Kaiser, and Vincent C. Marconi. 2017. Code Saturation Versus Meaning Saturation: How Many Interviews Are Enough? Qualitative Health Research 27, 4 (March 2017), 591–608. https://doi.org/10.1177/1049732316665344
- Holl et al. (2018) Jürgen Holl, Günter Kernbeiß, and Michael Wagner-Pinter. 2018. Das AMS-Arbeitsmarktchancen-Modell. Technical concept. Synthesis Forschung, Vienna.
- Jakesch et al. (2022) Maurice Jakesch, Zana Buçinca, Saleema Amershi, and Alexandra Olteanu. 2022. How Different Groups Prioritize Ethical Values for Responsible AI. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 310–323. https://doi.org/10.1145/3531146.3533097
- Kaun and Taranu (2020) Anne Kaun and Gabriela Taranu. 2020. Automating Society Report 2020 / Sweden. https://automatingsociety.algorithmwatch.org/report2020/sweden/
- Keil (2003) Frank C. Keil. 2003. Folkscience: coarse interpretations of a complex reality. Trends in Cognitive Sciences 7, 8 (Aug. 2003), 368–373. https://doi.org/10.1016/S1364-6613(03)00158-X
- Keil (2006) Frank C. Keil. 2006. Explanation and Understanding. Annual Review of Psychology 57, 1 (2006), 227–254. https://doi.org/10.1146/annurev.psych.57.102904.190100 arXiv:https://doi.org/10.1146/annurev.psych.57.102904.190100 PMID: 16318595.
- Keil (2021) Frank C. Keil. 2021. The Challenges and Benefits of Mechanistic Explanation in Folk Scientific Understanding. In Advances in experimental philosophy of science (paperback edition ed.), Daniel A. Wilkenfeld and Richard Samuels (Eds.). Bloomsbury Academic, London New York Oxford New Delhi Sydney, 41–56.
- Kelemen et al. (2013) Deborah Kelemen, Joshua Rottman, and Rebecca Seston. 2013. Professional physical scientists display tenacious teleological tendencies: Purpose-based reasoning as a cognitive default. Journal of Experimental Psychology: General 142, 4 (2013), 1074–1083. https://doi.org/10.1037/a0030399
- Kim et al. (2024) Minjung Kim, Saebyeol Kim, Jinwoo Kim, Tae-Jin Song, and Yuyoung Kim. 2024. Do stakeholder needs differ? - Designing stakeholder-tailored Explainable Artificial Intelligence (XAI) interfaces. International Journal of Human-Computer Studies 181 (2024), 103160. https://doi.org/10.1016/j.ijhcs.2023.103160
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. ”Help Me Help the AI”: Understanding How Explainability Can Support Human-AI Interaction. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages. https://doi.org/10.1145/3544548.3581001
- Kramer et al. (2018) Max F. Kramer, Jana Schaich Borg, Vincent Conitzer, and Walter Sinnott-Armstrong. 2018. When Do People Want AI to Make Decisions?. In Proceedings of the 2018 AAAI/ACM Conference on AI, Ethics, and Society (New Orleans, LA, USA) (AIES ’18). Association for Computing Machinery, New York, NY, USA, 204–209. https://doi.org/10.1145/3278721.3278752
- Kutsarova (2020) Viktoriya Kutsarova. 2020. Managing Alarming Situations with Mobile Crowdsensing Systems and Wearable Devices. KTH Royal Institute of Technology. School of Electrical Engineering and Computer Science. https://kth.diva-portal.org/smash/record.jsf?pid=diva2%3A1465193&dswid=-5365
- Kuziemski and Misuraca (2020) Maciej Kuziemski and Gianluca Misuraca. 2020. AI governance in the public sector: Three tales from the frontiers of automated decision-making in democratic settings. Telecommunications Policy 44, 6 (2020), 101976. https://doi.org/10.1016/j.telpol.2020.101976 Artificial intelligence, economy and society.
- Langer et al. (2021) Markus Langer, Daniel Oster, Timo Speith, Holger Hermanns, Lena Kästner, Eva Schmidt, Andreas Sesing, and Kevin Baum. 2021. What do we want from Explainable Artificial Intelligence (XAI)? – A stakeholder perspective on XAI and a conceptual model guiding interdisciplinary XAI research. Artificial Intelligence 296 (July 2021), 103473. https://doi.org/10.1016/j.artint.2021.103473
- Lau and Redlawsk (2001) Richard R. Lau and David P. Redlawsk. 2001. Advantages and Disadvantages of Cognitive Heuristics in Political Decision Making. American Journal of Political Science 45, 4 (Oct. 2001), 951. https://doi.org/10.2307/2669334
- Lee et al. (2019) Min Kyung Lee, Daniel Kusbit, Anson Kahng, Ji Tae Kim, Xinran Yuan, Allissa Chan, Daniel See, Ritesh Noothigattu, Siheon Lee, Alexandros Psomas, and Ariel D. Procaccia. 2019. WeBuildAI: Participatory Framework for Algorithmic Governance. Proceedings of the ACM on Human-Computer Interaction 3, CSCW (Nov. 2019), 1–35. https://doi.org/10.1145/3359283
- Liao et al. (2020) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3313831.3376590
- Liao et al. (2021) Q. Vera Liao, Milena Pribić, Jaesik Han, Sarah Miller, and Daby Sow. 2021. Question-Driven Design Process for Explainable AI User Experiences. arXiv:2104.03483 [cs.HC]
- Lim and Dey (2009) Brian Y. Lim and Anind K. Dey. 2009. Assessing Demand for Intelligibility in Context-Aware Applications. In Proceedings of the 11th International Conference on Ubiquitous Computing (Orlando, Florida, USA) (UbiComp ’09). Association for Computing Machinery, New York, NY, USA, 195–204. https://doi.org/10.1145/1620545.1620576
- Lima et al. (2023) Gabriel Lima, Nina Grgic-Hlaca, Jin Keun Jeong, and Meeyoung Cha. 2023. Who Should Pay When Machines Cause Harm? Laypeople’s Expectations of Legal Damages for Machine-Caused Harm. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 236–246. https://doi.org/10.1145/3593013.3593992
- Lombrozo and Gwynne (2014) Tania Lombrozo and Nicholas Z. Gwynne. 2014. Explanation and inference: mechanistic and functional explanations guide property generalization. Frontiers in Human Neuroscience 8 (Sept. 2014). https://doi.org/10.3389/fnhum.2014.00700
- Lombrozo and Wilkenfeld (2019) Tania Lombrozo and Daniel Wilkenfeld. 2019. Mechanistic versus Functional Understanding. In Varieties of Understanding. Oxford University Press, 209–230. https://doi.org/10.1093/oso/9780190860974.003.0011
- Long and Magerko (2020) Duri Long and Brian Magerko. 2020. What is AI Literacy? Competencies and Design Considerations. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–16. https://doi.org/10.1145/3313831.3376727
- Long et al. (2021) Duri Long, Aadarsh Padiyath, Anthony Teachey, and Brian Magerko. 2021. The Role of Collaboration, Creativity, and Embodiment in AI Learning Experiences. In Creativity and Cognition. ACM, Virtual Event Italy, 1–10. https://doi.org/10.1145/3450741.3465264
- Lopez (2019) Paola Lopez. 2019. Reinforcing Intersectional Inequality via the AMS Algorithm in Austria. (2019), 21.
- Lucaj et al. (2023) Laura Lucaj, Patrick van der Smagt, and Djalel Benbouzid. 2023. AI Regulation Is (Not) All You Need. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 1267–1279. https://doi.org/10.1145/3593013.3594079
- Lupia (2016) Arthur Lupia. 2016. Uninformed: why people know so little about politics and what we can do about it. Oxford University Press, New York.
- Madaio et al. (2020) Michael A. Madaio, Luke Stark, Jennifer Wortman Vaughan, and Hanna Wallach. 2020. Co-Designing Checklists to Understand Organizational Challenges and Opportunities around Fairness in AI. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–14. https://doi.org/10.1145/3313831.3376445
- Miller (2019) Tim Miller. 2019. Explanation in artificial intelligence: Insights from the social sciences. Artificial Intelligence 267 (Feb. 2019), 1–38. https://doi.org/10.1016/j.artint.2018.07.007
- Miller (2023) Tim Miller. 2023. Explainable AI is Dead, Long Live Explainable AI! Hypothesis-Driven Decision Support Using Evaluative AI. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 333–342. https://doi.org/10.1145/3593013.3594001
- Mueller et al. (2019) Shane T. Mueller, Robert R. Hoffman, William Clancey, Abigail Emrey, and Gary Klein. 2019. Explanation in Human-AI Systems: A Literature Meta-Review, Synopsis of Key Ideas and Publications, and Bibliography for Explainable AI. arXiv:1902.01876 [cs.AI]
- Nannini et al. (2023) Luca Nannini, Agathe Balayn, and Adam Leon Smith. 2023. Explainability in AI Policies: A Critical Review of Communications, Reports, Regulations, and Standards in the EU, US, and UK. In 2023 ACM Conference on Fairness, Accountability, and Transparency. ACM, Chicago IL USA, 1198–1212. https://doi.org/10.1145/3593013.3594074
- Ng et al. (2021) Davy Tsz Kit Ng, Jac Ka Lok Leung, Kai Wah Samuel Chu, and Maggie Shen Qiao. 2021. AI Literacy: Definition, Teaching, Evaluation and Ethical Issues. Proceedings of the Association for Information Science and Technology 58, 1 (Oct. 2021), 504–509. https://doi.org/10.1002/pra2.487
- Niklas et al. (2015) Jedrzej Niklas, Karolina Sztandar-Sztanderska, and Katarzyna Szymielewicz. 2015. Profiling the Unemployed in Poland: Social and Political Implications of Algorithmic Decision Making. (10 2015).
- on Artificial Intelligence (2019) High-Level Expert Group on Artificial Intelligence. 2019. Ethics guidelines for trustworthy AI. (April 2019). https://digital-strategy.ec.europa.eu/en/library/ethics-guidelines-trustworthy-ai
- Páez (2019) Andrés Páez. 2019. The Pragmatic Turn in Explainable Artificial Intelligence (XAI). Minds and Machines 29, 3 (01 Sep 2019), 441–459. https://doi.org/10.1007/s11023-019-09502-w
- Panigutti et al. (2023) Cecilia Panigutti, Ronan Hamon, Isabelle Hupont, David Fernandez Llorca, Delia Fano Yela, Henrik Junklewitz, Salvatore Scalzo, Gabriele Mazzini, Ignacio Sanchez, Josep Soler Garrido, and Emilia Gomez. 2023. The Role of Explainable AI in the Context of the AI Act. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 1139–1150. https://doi.org/10.1145/3593013.3594069
- Rozenblit and Keil (2002) Leonid Rozenblit and Frank Keil. 2002. The misunderstood limits of folk science: an illusion of explanatory depth. Cognitive Science 26, 5 (2002), 521–562. https://doi.org/10.1207/s15516709cog2605_1 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1207/s15516709cog2605_1
- Rugg and Petre (2007) Gordon Rugg and Marian Petre. 2007. A gentle guide to research methods. Open University Press.
- Sahlen et al. (2018) Malin Sahlen, Edvin Alam, and Paulo Silva. 2018. Ny teknik i välfärden - vår tids välfärdskliv (New Technologies in Welfare – The Welfare Leap of Our Times). Technical Report. Svenskt Näringsliv. https://www.svensktnaringsliv.se/sakomraden/valfard-och-offentlig-sektor/ny-teknik-i-valfarden-var-tids-valfardskliv_1003696.html
- Sato et al. (2019) Brian K. Sato, Cynthia F. C. Hill, and Stanley M. Lo. 2019. Testing the test: Are exams measuring understanding? Biochemistry and Molecular Biology Education 47, 3 (May 2019), 296–302. https://doi.org/10.1002/bmb.21231
- Schmude et al. (2023a) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2023a. On the Impact of Explanations on Understanding of Algorithmic Decision-Making. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 959–970. https://doi.org/10.1145/3593013.3594054
- Schmude et al. (2023b) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2023b. Applying Interdisciplinary Frameworks to Understand Algorithmic Decision-Making. https://doi.org/10.48550/arXiv.2305.16700 arXiv:2305.16700 [cs.HC]
- Schoeffer et al. (2022) Jakob Schoeffer, Niklas Kuehl, and Yvette Machowski. 2022. “There Is Not Enough Information”: On the Effects of Explanations on Perceptions of Informational Fairness and Trustworthiness in Automated Decision-Making. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1616–1628. https://doi.org/10.1145/3531146.3533218
- Scott et al. (2022) Kristen M. Scott, Sonja Mei Wang, Milagros Miceli, Pieter Delobelle, Karolina Sztandar-Sztanderska, and Bettina Berendt. 2022. Algorithmic Tools in Public Employment Services: Towards a Jobseeker-Centric Perspective. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 2138–2148. https://doi.org/10.1145/3531146.3534631
- Shen et al. (2020) Hong Shen, Haojian Jin, Ángel Alexander Cabrera, Adam Perer, Haiyi Zhu, and Jason I. Hong. 2020. Designing Alternative Representations of Confusion Matrices to Support Non-Expert Public Understanding of Algorithm Performance. Proceedings of the ACM on Human-Computer Interaction 4, CSCW2 (Oct. 2020), 1–22. https://doi.org/10.1145/3415224
- Shin (2021) Donghee Shin. 2021. The effects of explainability and causability on perception, trust, and acceptance: Implications for explainable AI. International Journal of Human-Computer Studies 146 (2021), 102551. https://doi.org/10.1016/j.ijhcs.2020.102551
- Shulner-Tal et al. (2022) Avital Shulner-Tal, Tsvi Kuflik, and Doron Kliger. 2022. Enhancing Fairness Perception – Towards Human-Centred AI and Personalized Explanations Understanding the Factors Influencing Laypeople’s Fairness Perceptions of Algorithmic Decisions. International Journal of Human–Computer Interaction (July 2022), 1–28. https://doi.org/10.1080/10447318.2022.2095705
- Sovrano et al. (2022) Francesco Sovrano, Salvatore Sapienza, Monica Palmirani, and Fabio Vitali. 2022. Metrics, Explainability and the European AI Act Proposal. J 5, 1 (2022), 126–138. https://doi.org/10.3390/j5010010
- Szymanski et al. (2021) Maxwell Szymanski, Martijn Millecamp, and Katrien Verbert. 2021. Visual, textual or hybrid: the effect of user expertise on different explanations. In 26th International Conference on Intelligent User Interfaces. ACM, College Station TX USA, 109–119. https://doi.org/10.1145/3397481.3450662
- Thiebes et al. (2021) Scott Thiebes, Sebastian Lins, and Ali Sunyaev. 2021. Trustworthy Artificial Intelligence. Electronic Markets 31, 2 (June 2021), 447–464. https://doi.org/10.1007/s12525-020-00441-4
- Veale et al. (2018) Michael Veale, Max Van Kleek, and Reuben Binns. 2018. Fairness and Accountability Design Needs for Algorithmic Support in High-Stakes Public Sector Decision-Making. Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (April 2018), 1–14. https://doi.org/10.1145/3173574.3174014
- Veale and Zuiderveen Borgesius (2021) Michael Veale and Frederik Zuiderveen Borgesius. 2021. Demystifying the Draft EU Artificial Intelligence Act. preprint. SocArXiv. https://doi.org/10.31235/osf.io/38p5f
- Vredenburgh (2022) Kate Vredenburgh. 2022. The Right to Explanation*. Journal of Political Philosophy 30, 2 (June 2022), 209–229. https://doi.org/10.1111/jopp.12262
- Wang and Yin (2021) Xinru Wang and Ming Yin. 2021. Are Explanations Helpful? A Comparative Study of the Effects of Explanations in AI-Assisted Decision-Making. In 26th International Conference on Intelligent User Interfaces (IUI ’21). Association for Computing Machinery, New York, NY, USA, 318–328. https://doi.org/10.1145/3397481.3450650
- Wiggins and McTighe (2005) Grant P. Wiggins and Jay McTighe. 2005. Understanding by design (expanded 2nd ed ed.). Association for Supervision and Curriculum Development, Alexandria, VA. OCLC: 60756429.
- Woodruff et al. (2018) Allison Woodruff, Sarah E. Fox, Steven Rousso-Schindler, and Jeffrey Warshaw. 2018. A Qualitative Exploration of Perceptions of Algorithmic Fairness. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. ACM, Montreal QC Canada, 1–14. https://doi.org/10.1145/3173574.3174230
- Zagzebski (2019) Linda Zagzebski. 2019. Toward a Theory of Understanding. In Varieties of Understanding. Oxford University Press, 123–136. https://doi.org/10.1093/oso/9780190860974.003.0007
- Zerilli (2022) John Zerilli. 2022. Explaining Machine Learning Decisions. Philosophy of Science 89, 1 (2022), 1–19. https://doi.org/10.1017/psa.2021.13
- Züger and Asghari (2023) Theresa Züger and Hadi Asghari. 2023. AI for the public. How public interest theory shifts the discourse on AI. AI & SOCIETY 38, 2 (April 2023), 815–828. https://doi.org/10.1007/s00146-022-01480-5